# Quantifying sunspot group nesting with density-based unsupervised clustering

### v2

@@ 3.5 Inter-nest Length Scale
$L_{\rm typ}=(A/N_{\rm sig})^{1/2},$ (9)
where $A=360^{\circ}\times 5^{\circ}$ is the total surface area of the latitude band, and $N_{\rm sig}$ is the number of statistically significant nests identified in that window. This metric represents the mean separation obtained by uniformly distributing $N_{\rm sig}$ nests across the available longitude range. We convert this angular separation to physical distance (Mm) using a reference latitude of $15^{\circ}$ ($\cos(15^{\circ})\simeq 0.966$), representative of the mid-latitude active region belt where most nesting occurs.
@@
Figure 10 shows the inter-nest length scale as a function of total SG area (a proxy for solar activity level) for both RGO (1874–1976) and KMAS (1954–2025) latitude-time windows. Both datasets exhibit a strong negative correlation between $L_{\rm typ}$ and activity level, with Pearson coefficients 0.647-0.647 ($p\sim 10^{-15}$) for RGO and 0.682-0.682 ($p\sim 10^{-13}$) for KMAS. At low activity levels (sparse emergence), $L_{\rm typ}$ reaches 200-500 Mm, while at high activity levels it converges to $\sim$60-100 Mm, comparable to typical SG elongations in the E-W direction (3–10, corresponding to $\sim$35–117 Mm at $15^{\circ}$ latitude). The mean value of $L_{\rm typ}$ is approximately 125 Mm for both datasets.
@@ 4 Discussion
The inter-nest length scale $L_{\rm typ}$ (Sect. 3.5) represents a phase-space density measure integrating over full cycle durations and uniform spatial distribution across $360^{\circ}\times 5^{\circ}$ bands. Consequently, $L_{\rm typ}$ does not reflect actual local spacing between neighbouring nests within active zones, where emergence concentrates preferentially (Gaizauskas et al., 1983). Nevertheless, the robust anti-correlation with activity ($r\approx-0.65$, $p<10^{-13}$) reveals a significant pattern: during solar maximum, nests pack into denser configurations, approaching typical SG elongations. This convergence may reflect hierarchical magnetic organisation, with compact nests clustering into larger composite structures spanning up to $55^{\circ}$ longitude (Brouwer and Zwaan, 1990). Refined analysis focusing on active emergence periods rather than cycle-integrated densities would better characterise local inter-nest spacing but lies beyond the scope of the present study.

### v3

@@ 3.5 Inter-nest Length Scale
$\lambda_{\rm typ}=(A/N_{\rm sig})^{1/2},$ (9)
where $A=360^{\circ}\times 5^{\circ}$ is the angular area of the latitude band, and $N_{\rm sig}$ is the number of statistically significant nests identified in that window. This metric represents the mean angular separation obtained by uniformly distributing $N_{\rm sig}$ nests across the available longitude range. We convert this to a physical length scale $L_{\rm typ}$ (Mm) using the central latitude of each band.
@@
Figure 10 shows the inter-nest length scale as a function of total SG area (a proxy for solar activity level) for both RGO (1874–1976) and KMAS (1954–2025) latitude-time windows. Both datasets exhibit a strong negative correlation between $L_{\rm typ}$ and activity level, with Pearson coefficients 0.641-0.641 ($p\sim 10^{-16}$) for RGO and 0.700-0.700 ($p\sim 10^{-15}$) for KMAS. At low activity levels (sparse emergence), $L_{\rm typ}$ reaches 200–500 Mm, while at high activity levels it converges to $\sim$60–100 Mm, comparable to typical SG elongations in the E-W direction (3–10, corresponding to $\sim$35–117 Mm at $15^{\circ}$ latitude). The mean value of $L_{\rm typ}$ is approximately 120 Mm for both datasets.
@@ 4 Discussion
The inter-nest length scale $L_{\rm typ}$ (Sect. 3.5) represents a phase-space density measure integrating over full cycle durations and uniform spatial distribution across $360^{\circ}\times 5^{\circ}$ bands. Consequently, $L_{\rm typ}$ does not reflect actual local spacing between neighbouring nests within active zones, where emergence concentrates preferentially (Gaizauskas et al., 1983). Nevertheless, the robust anti-correlation with activity (Fig. 10) reveals a significant pattern: during solar maximum, nests pack into denser configurations, approaching typical SG elongations. This convergence may reflect hierarchical magnetic organisation, with compact nests clustering into larger composite structures spanning up to $55^{\circ}$ longitude (Brouwer and Zwaan, 1990). Refined analysis focusing on active emergence periods rather than cycle-integrated densities would better characterise local inter-nest spacing but lies beyond the scope of the present study.